### v6

# Involution factorizations of Ewens random permutations

Abstract
An involution is a bijection that is its own inverse. Given a permutation of let denote the number of ways can be expressed as a composition of two involutions of We prove that the statistic is asymptotically lognormal when the symmetric groups are each equipped with Ewens Sampling Formula probability measures of some fixed positive parameter This paper strengthens and generalizes previously determined results about the limiting distribution of for uniform random permutations, i.e. the specific case of . We also investigate the first two moments of itself, detailing the phase transition in asymptotic behavior at and provide a functional refinement and a demonstrably optimal convergence rate for the log-Gaussian limit law.
keywords:
asymptotic lognormality, Erdős-Turán law, Ewens Sampling Formula, Feller coupling, involutions, Mellin transform, saddle-point method1 Introduction
An involution is a bijection that is its own inverse. The involutions in the symmetric group are precisely those permutations whose cycles are all of length 1 or 2. This paper is concerned with decomposing random permutations into two involutive factors and the probabilistic analysis thereof. Recent interest in involution factorizations in can be traced to Roberts and Vivaldi (2009) who proposed that the dynamical behavior of reversible rational maps over a finite field can be modeled by compositions of two random involutions with a prescribed number of fixed points. Unrestricted products of two involutions, however, are of singular interest as well. For instance, Petersen and Tenner (2013) outline hypothetical links between enumerating involution factorizations, rim-hook tableaux, and the irreducible characters of that are still in need of further clarification. Other applications have surfaced in comparative genomics (e.g. Bhatia et al. (2015), Chindelevitch and Meidanis (2017)) and the analysis of algorithms (e.g. Kawahara et al. (2019) and Yang et al. (2013)).
If is a permutation of let denote the number of function compositional factorizations of into the product of two involutions of that is, the number of ordered pairs of involutions in such that It is known that
where is the falling factorial denotes the number of cycles of length (or, more succinctly, -cycles) that possesses, and is the product of its cycle lengths. To the author’s knowledge, the earliest proofs of (1.1) are independently attributed to Lugo (2009b) and Petersen and Tenner (2013). Both of their derivations rely on identifying involutions with partial matchings and, consequently, the composition of two involutions with dichromatic edge-colored graphs in which every vertex is incident to exactly two edges, one edge per color. A derivation of (1.1) that merely exploits the canonical group action of on appears in Burnette and Schmutz (2016).
Even without (1.1), one notices that is a class function by virtue of the conjugacy invariance of the set of involutions. It is therefore natural to wonder what the distribution of is for permutations drawn from conjugation-invariant probability measures on where cycles are multiplicatively weighted. The archetypal example of such a measure is the Ewens Sampling Formula with parameter hereafter abbreviated Read the survey article by Crane (2016) for a lucid and concise synopsis of the literature on
For each positive integer the Ewens probability measure is a deformation of the uniform measure on obtained by incorporating the following sampling bias:
where is the rising factorial and is the total number of cycles that has. (Note, in particular, that ) More specifically, the probability that a permutation of chosen according to has cycle index is
where denotes the indicator function. We shall also refer to random permutations chosen according ot as -weighted permutations, as is also done in Lugo (2009a).
A pivotal feature of is that the joint distribution of the cycle counts converges to that of independent Poisson random variables of respective intensity Now consider a sequence of mutually independent random variables with for and define
Note that, as
whereas
Markov’s inequality thus dictates that
One would then suspect that something similar holds for and in assuming both are plausibly well-approximated by their Poissonized counterparts.
In this paper, we will establish both central limit and functional limit theorems for over Ewens random permutations. This will be accomplished by generalizing the key results of Burnette and Schmutz (2016). It was particularly shown in Burnette and Schmutz (2016) that is asymptotically lognormal when each is endowed with the uniform probability measure. The proof of this hinged on the observation that, for asymptotically almost all permutations the statistic is readily controlled by the simpler statistic together with the fact that is itself an asymptotically lognormal random variable over uniform random permutations. As we shall soon witness, these facts remain true in the setting of The polynomial factors of (1.1) ony influence the remainder terms in the weak limit theorems stated in Section 5.
On the other hand, the distribution of itself, when left untouched by any rescaling or standardization, is heavily right-skewed. Practically all of its range is concentrated far below its mean. To demonstrate this, we will derive the leading-term asymptotics of the first and second moments of for general -weighted permutations. The latter will highlight certain discrepancies in the variability of for versus
We remark that the limiting distribution of falls within the scope of the general theory of additive functionals on the cycle multiplicities of random permutations (see e.g. Manstavichyus (1996), Babu and Manstavichyus (2002)). The original contribution of the present paper, however, is the singularity analysis of the generating functions that are used to concretely describe the distribution of Moreover, all of the fine moment asymptotics that can be found throughout Sections 2 through 5 have not appeared in the literature before.
Some classical results and techniques from asymptotic analysis are cited in a piecemeal fashion throughout this paper. As is common for random permutation and integer partition statistics, we will encounter generating functions that demand a careful examination by way of the saddle-point method and Mellin transforms. Technical preliminaries are briefly discussed in the appendix, Section 6, at the end.
2 The Expected and Typical Values of
Let denote the probability measure on corresponding to For the sake of notational brevity, we will henceforth omit the in and associated operators.
Theorem 2.1.
The average number of involution factorizations for a permutation chosen according to the Ewens measure is
Proof.
For each set Consider the power series
By the law of total expectation,
where means that and As pointed out in Theorem 2.5 of Burnette and Schmutz (2016),
where is the probabilists’ Hermite polynomial Thus
Invoking the exponential generating function (an assortment of derivations are available in Kim et al. (2012), Kim (2012), Kim and Mansour (2014)), we see that
Hence
Applying Corollary 6.2 with and together with the Stirling’s approximations and completes the proof. ∎
We can further adapt the function constructed in the proof of Theorem 2.1 to compute the average number of involution factorizations for permutations with a fixed number of cycles. Indeed, if we set
then
is the bivariate ordinary generating function for the expected value of over uniform random permutations, counted according to the size of the ground set (indicated by the variable ) and the total number of cycles (indicated by the variable ). The diagonal of is given by
which is the exponential generating function for the class of involutions. This makes sense, as the only permutation of with cycles is the identity permutation, which is precisely the square of every involution in By expanding the exponential series in (2.11), we can work towards retrieving the conditional expectation of for permutations with cycles by using the “vertical” generating function
For example, and so for This too makes sense as the probability of picking an -cycle is and there are exactly ways to write an -cycle as a product of two involutions.
Theorem 2.2.
For each positive integer
as
Proof.
Because is a sufficient statistic for the underlying parameter which means that the conditional distribution of does not depend on (see section 2.2 of Crane (2016)), the conditional expectation is constant across the support of Therefore
where are the unsigned Stirling numbers of the first kind. Based on the Flajolet-Odlyzko transfer theorem (see Flajolet and Odlyzko (1990))
we find that
Evoke the following crude simplification of Wilf’s asymptotic (Wilf (1993))
to finish the calculation. ∎
Most of the contribution to the mean comes from a small subset of permutations that admit an exceptionally large number of involution factorizations. Scrupulous glances at (1.1) should convince the reader that giving a permutation numerous occurrences of a distinct long cycle length or an exorbitant number of short cycles exaggerates the magnitude of But it is rare for a permutation to have such a cycle profile. The typical permutation has much fewer involution factorizations. Direct combinatorial arguments can verify this for uniform random permutations, which have roughly involution factorizations on average whereas the median is which we will show in Theorem 2.4 and its corollary. (In fact, it is closer to as shown in Burnette and Schmutz (2016).) For this, we use the following lemma.
Lemma 2.3.
For each set Then is attained at a permutation with the maximally obtainable amount of shortest cycles.
For a cursory explanation as to why, note that if and is a partition of with then any permutation obtained by splitting up each -cycle of into -cycles for generally enlarges the number of involution factorizations. (Cf the first two paragraphs in the proof of Theorem 16 in Schneider and Sills (2020) which discuss how this strategy optimizes the product of the parts of a partition.)
In what follows, we let and be integers such that and where is the triangular number. Let be the set of all permutations for which for all and for all Observe that the condition forces each permutation in to have at least one cycle longer than Lemma III of Erdös and Turán (1965) states that
We can adapt this result to bound the median number of involution factorizations in uniform random permutations.
Theorem 2.4.
If denotes the harmonic number and
then at least permutations satisfy the inequality
Proof.
Let denote the number of distinct cycle lengths present in the cycle decomposition of and let be the event that It is folklore that and are the mean and standard deviation, respectively, of over uniform random permutations of This together with (2.18), the union bound, and Chebyshev’s inequality tell us that
Now let be a permutation for which attains its maximum value over By Lemma 2.3, if we set then for and for up to an additional larger integers Hence,
Setting in Theorem 2.4 and employing the Stirling bound from Robbins (1955)
along with the elementary bounds and begets the following.
Corollary 2.5.
If but slowly enough so that then all but permutations satisfy the inequality
Goncharov’s limit law for over (Theorem 3.5, Arratia et al. (2000)) and the upcoming Lemma 4.2 indicate that, assuming the same notations of the previous proof, asymptotically almost all -weighted permutations satisfy the analogous inequality
However, we can conclude something even stronger. Based on Theorem 4.4, if is a -weighted permutation of and then w.h.p.,
3 A Close Look at the Second Moment of
Ultimately, the asymptotic lognormality of for -weighted permutations prevents, say, the method of moments from succeeding. (See, for instance, Example 30.2 on pg. 389 of Billingsley (1995).) Ignoring that, we can still observe that the second moment of is quite obstreperous; its growth rate vastly outpaces the first moment, and the extent to which it does depends on whether .
For a hint at this disparity, note that the Ewens Sampling Formula favors permutations with many cycles whenever This inflates the likelihood of drawing a permutation with a lot of short cycles, which is when is at its biggest. So grows drastically faster than it does in the case and falls under an entirely different asymptotic regime.
The ensuing analysis is included here not only to contextualize the sheer futility of the method of moments in this situation but also to help cast light on the statistical spread of for -weighted permutations and how acts as a “phase transition point” for the variance. Reprising the notation presented in the proof of Theorem 2.1, we consider the generating function
We now recall the probabilists’ version of the Mehler kernel (see Kibble (1945))
to observe that
Theorem 3.1.
If then
Proof.
Note that has a dominant singularity at and the function given by
is regular in the disc Corollary 6.2 can thus be applied to get
where
Hence
Here we see that However, if then has a radius of convergence of 1, but behaves too chaotically around the unit circle to be analytically continued beyond it. In this case, we appeal to the exp-log reorganization of (3.5):
From this, we can attain the following upper bound on for when
where is the polylogarithm function of order (Apply the arithmetic mean-harmonic mean inequality on the terms of the second summation to transition from (3.15) to (3.16).)
Now in setting the function
is a harmonic sum well-suited for a Mellin transform analysis. Consult section 6.2 of the appendix to deduce that the Mellin transform of is
Note that when treated as a function of and with held constant, is analytic for because converges absolutely for all when If then the polylogarithm is simply the Riemann zeta function which has a simple pole at We therefore consider the cases and separately.
Theorem 3.2.
If then, as
Proof.
The Mellin transform has simple poles at Using the standard transfer rule for the transform, we ascertain that
which, because of the Laurent series expansions of the powers of at and the fact that (3.18) extends to corresponds to
from within the unit disc. We can now conclude that
provided with
We claim next that To see this, consider the positively oriented circle and split into the two contours
and By Cauchy’s integral formula,
The length of is This together with (3.20) shows that the first integral over is
On the other hand, estimate (3.17) applies over and so the second integral in (3.22) is
for some Therefore
It is now a sophomore’s homework exercise to use Theorem 6.1 on especially considering that the required singular expansion is already accounted for in (3.18). The error term in (3.29) is smaller than the leading term asymptotic of by a decaying factor of In the end, we find that
and so the result follows again from the fact that ∎
Theorem 3.3.
If then
Proof.
Here, is precisely the exponential generating function of and the Mellin transform becomes
which has simple poles at and a double pole at Hence
which corresponds to
from within the unit disc. Therefore
provided and
At this point, the outstanding justifications are practically the same as they were for the case. It is tedious, but requires nothing new from earlier to figure out that
4 A Central Limit Theorem for
In this section, we emulate the proof of Theorem 4.3 in Burnette and Schmutz (2016), except now each is equipped with the ESF probability measure with fixed positive . We proceed in three steps. First, is well-approximated by for permutations in which only -cycles with under a certain threshold to be stipulated a posteriori, can occur more than once but still not more than times. Call the set of all such permutations The second step is to show that as Lastly, the previous two steps are combined to compare the distribution of to the distribution of
There are two notable improvements here over the presentation of Burnette (2017) and Burnette and Schmutz (2016). Our present proof of the aforementioned second step is cleaner and considerably streamlined. Furthermore, a concrete error term for the approximate Gaussian law will be added in the next section. Perhaps unsurprisingly, the error term matches the convergence rate of the Erdős-Turán limit law for the order of a random permutation that was first determined in Barbour and Tavaré (1994).
Lemma 4.1.
(Burnette and Schmutz (2016)) Suppose Then there is a constant not dependent on nor such that
Lemma 4.1 is entirely deterministic and thus needs not be revisited. The next lemma generalizes Lemma III in Erdös and Turán (1965).
Lemma 4.2.
If then
Proof.
Due to (1.3) and the exponential formula for labeled combinatorial structures,
The product in (4.1) can be written as where
Equating the corresponding coefficients reveals that
If we replace every instance of the term by in each factor of we obtain an entire function whose coefficients are all positive and majorize the absolute values of the corresponding coefficients of Hence
For sufficiently large we have
for some constant and
for some constant Therefore, as
We now cite the asymptotic lognormality of for -weighted permutations.
Lemma 4.3.
Erdös and Turán (1967) were the first to confirm the asymptotic lognormality of for uniform random permutations. Progressively sharper versions of the Erdős-Turán limit law for have emerged on multiple occasions, e.g. Bovey (1980), DeLaurentis and Pittel (1985), Manstavicius (1998), Nicolas (1985), Zakharovas (2002). Proofs of Lemma 4.3 itself appear in the aggregate works of Arratia, Barbour, and Tavaré throughout Arratia et al. (2000), Arratia et al. (2003), Arratia and Tavaré (1992), Barbour and Tavaré (1994).
Theorem 4.4.
For all real
Proof.
Because for all an upper bound is immediate:
For a lower bound, we capitalize on the uniform continuity of We also invoke the conditional bound from Lemma 4.1, where will now be chosen to our liking. Due to Lemma 4.2, this bound holds with probability
The remainder of the proof is virtually identical to that of Theorem 4.3 in Burnette and Schmutz (2016), which is included here for the sake of completeness. Let be a fixed but arbitrarily small positive number. We can choose so that whenever If we choose then we have Therefore we can choose so that, for all , we have It follows that, uniformly in
Yet was arbitrary, and so ∎
5 A Functional Refinement and Convergence Rate
A stochastic process variant of Theorem 4.4 can also be formulated. Let be distributed according to and be the random element of the Skorokhod space of càdlàg functions on defined by
Theorem 5.1.
It is possible to construct and a standard Brownian motion on the same probability space in such a way that
The Feller coupling, which expresses the Ewens Sampling Formula in terms of the spacings between successes in a sequence of independent Bernoulli trials, offers the ideal terrain for this task. Arratia et al. (2016) wrote a brief expository note about the Feller coupling to act as a companion to Crane (2016). To summarize, consider a sequence of independent Bernoulli random variables in which Say that a -spacing occurs in a given binary sequence , starting at position and ending at position if and for If for all positive integers and we define
it can be shown that the distribution of is exactly the same as it is in If we further define where
then as due to the Borel-Cantelli lemma, and so is equivalent to the Poisson point process to which itself converges.
We now assemble the pieces needed to prove Theorem 5.1. Our plan is to show that within the Feller coupling, is “close enough” to which in turn is “close enough” to which is known to conform to a Wiener process. To this end, we raise two lemmas. The first is a trivial consequence of the preceding definitions.
Lemma 5.2.
For all positive integers and
where and
Lemma 5.3.
If and are two vectors such that for then
for all where denotes the usual -norm.
Proof.
Observe that
where we have set
Note that for all positive integers and ∎
Lemma 5.4.
Proof.
The remainder of the proof of Theorem 5.1 is immediate in light of the following.
Theorem 5.5.
(Barbour and Tavaré (1994)) Suppose for all we define
Then converges weakly in to a standard Brownian motion in the sense that
We now articulate the proximity of the quantities and with the goal of appending a convergence rate to Theorem 4.4. Let us introduce the two standardized random variables
The following generalization of Lemma 2.5 in Barbour and Tavaré (1994) will help us show that the distributions of and are almost the same.
Lemma 5.6.
Let and be random variables. Suppose that
where is a Lipschitz function over Then, for any
where the constant depends only on
Theorem 5.7.
If is distributed according to then
Proof.
We begin by applying Lemma 5.6 with and By Lemma 4.3, there exists a constant such that So if we take and then applying a generic Chernoff bound yields
Leverage the fact that for at most one integer to see that
Yet since the are pairwise independent and we can interchange the product and expectation operator and then compute the factorial moments to get
Therefore and so
According to a result of Zakharovas (2002),
Putting into (5.6),
and observing that
yields the following.
Theorem 5.8.
If is distributed according to then
Zacharovas’s result on the optimality of the convergence rate for the Erdős-Turán limit law is constructive and can be mimicked for -weighted permutations in general. For this reason, the rate of convergence stated in Theorem 5.7 is the best possible. The proof is excluded here, but all of the relevant details can be found in Zakharovas (2002).
Although we decline from doing so here, we remark that it is straightforward to compute explicit bounds for the constants implied by the error terms in Theorems 5.1 and 5.7. The estimates needed to measure the distance between the distributions of and already appear in Barbour and Tavaré (1994). From there, a Berry-Esséen inequality will disclose how quickly tends to its Gaussian limit law. The constants implied by (5.11) and (5.12) require more delicate asymptotic evaluation, but can be safely ignored since the dominant error term for Theorem 5.1 resides in Theorem 5.5.
Concluding Remarks
The nonuniformity of makes it evident that the composition of two uniformly random involutions is not a uniformly random permutation. Furthermore, if we let denote the total number of involutions in then
(Refer to pg. 583, Example VIII.9, Flajolet and Sedgewick (2009).) Therefore, using pairs of random involutions to generate random permutations introduces a significant sampling bias.
Like for a composition of two random involutions is likelier to have a lot of cycles. However, the distribution of looks quite different in this model. The factors of the exponential generating function in (2.6) suggests that converges in law to the distribution of where and are independent Poisson random variables of mean 1 and respectively. Lugo (2009b) proved as much in an unpublished manuscript. The author intends to expand on these and other observations in a followup paper.
Acknowledgements.
This paper refines various elements of the author’s doctoral thesis. The author wishes to thank his advisor, Eric Schmutz, for his extensive guidance. The author also expresses his gratitude to Dirk Zeindler for first suggesting to him that the results of Burnette and Schmutz (2016) can be generalized to the Ewens Sampling Formula, to Hsien-Kuei Hwang for helpful discussions on how to tackle the variance of and to Eugenijus Manstavičius for reaching out and sharing some insightful background on this topic.6 Appendix
This appendix compiles some technical prerequisites which are used sporadically throughout the manuscript.
6.1 Wright’s Expansions
The following consequence of the saddle-point method applied to functions with exponential singularities is often helpful in unraveling the asymptotic development of various random permutation statistics.
Theorem 6.1.
(Wright (1949)) Suppose
where is regular in the unit disc,
are real, are complex, and If in addition
where and the and the are all calculable in terms of the and then
where
With one exception in section 3, we solely need this simpler instance of Theorem 6.1.
Corollary 6.2.
(Wright (1932)) The leading-term asymptotic for
where is a complex number, is regular in the unit disc, and is a nonzero real number, is given by
6.2 The Mellin Transform
To conclude this section, we review some pertinent facts about the Mellin transform as a means of extracting asymptotic estimates. Given a locally integrable function defined over the Mellin transform of is the complex-variable function
The transform is also occasionally denoted by or If as and as then is an analytic function on the “fundamental strip” and is often continuable to a meromorphic function on the whole of
Notably, the Mellin transform is a linear operator that satisfies the rescaling rule
for any This together with the linearity of yields the “harmonic sum rule”
provided that the integral and summation are interchangeable. Sums of the form are known as harmonic sums with base function amplitudes and frequencies In other words, the Mellin transform of a harmonic sum factors into the product of the Mellin transform of the base function and a generalized Dirichlet series associated with the corresponding sequence of amplitudes and frequencies. For instance, we can obtain the useful Mellin pairs
and
Akin to many other valuable integral transforms, the Mellin transform is essentially involutive. If is continuous in an interval containing then can be recovered from via the inversion formula
where the abscissa must be chosen from the fundamental strip of If the continuation of decays quickly enough as then the inverse Mellin integral and Cauchy’s residue theorem can be combined to derive an asymptotic series for as We have the following transfer rule in particular: a pole of order at with
translates to the term in the singular expansion of near (The specific requirements for this transfer rule are all fulfilled by the special functions appearing in this manuscript, as the gamma function decays rapidly along vertical lines in the complex plane whereas polylogarithms undergo only moderate growth.) Detailed explanations along with further uses of the Mellin transform towards solving asymptotic enumeration problems can be found in Flajolet et al. (1995) and Flajolet and Sedgewick (2009).